OpenStudy (anonymous):

solve for theta [0,2pi] sin^4(theta)+4cos^2(theta)=1

OpenStudy (anonymous):

Hint :\[\sin^4 \theta+\cos^4 \theta=(\sin^2 \theta+\cos^2 \theta)^2-2\sin^2 \theta \cos^2 \theta\]

OpenStudy (anonymous):

wait there is a 4 there sorry

OpenStudy (anonymous):

maybe write\[\sin^4 \theta=(1-\cos^2 \theta)^2\]

OpenStudy (anonymous):

It is obvious that \[ \theta =\pm \frac\pi 2 \] are solutions

OpenStudy (anonymous):

so pi/2 and 3pi/2

OpenStudy (anonymous):

are those the only solutions?

OpenStudy (anonymous):

I think there are some more. Stay tuned.

OpenStudy (anonymous):

Also we can show that the equation is equivalent to \[ \cos ^2(\theta ) \left(5 \cos ^2(\theta )+2\right)=0 \]

OpenStudy (anonymous):

so, the only answers are pi/2 and 3pi/2

OpenStudy (anonymous):

No the other ones are in your interval too.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Wait a minute. I am examining something.

OpenStudy (anonymous):

Only pi/2 and 3pi/2

OpenStudy (anonymous):

thanks. :)

OpenStudy (anonymous):

Because \(5 \cos ^2(\theta )+2\ >0\)

OpenStudy (anonymous):

yw